Circular No. 6142

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

G299.2–2.9 J.-U. Busser, R. J. Egger, and B. Aschenbach, Max-Planck-Institut

für Extraterrestrische Physik, Garching, report: “A new supernova remnant, G299.2–2.9, centered at α = 12^{h}15^{m}11^{s}.6, δ = −65^{o}30′04′′ (equinox 2000.0), has been discovered in the ROSAT All-Sky Survey. G299.2–2.9 is nearly circular with a diameter of 15′ and has a centrally-filled shape. The European Southern Observatory’s red survey plate No. 92 shows nebulous filaments around the rim of the source, which also shows up as a ring in the 60- to 100-μm IRAS ratio map. No radio counterpart has been found in available radio catalogues. A thermal plasma model best fit reveals a temperature of ∼ 8 × 10^{6} K and a column density of N_{H} ∼ 3 × 10^{20} cm^{−2}. With a mean interstellar H I density of 0.2 cm^{−3}, the distance to G299.2– 2.9 is ∼ 500 pc. The observed count rate in the ROSAT band (0.1–2.4 keV) is 3.3 counts/s and the equivalent unabsorbed energy flux is 2.3 × 10^{−11} erg cm^{−2} s^{−1}. At the distance of 500 pc, the derived ambient density is 0.2 cm^{−3} and the x-ray radiating mass is < 0.1 M_{⊙}, which suggests that the remnant is comparably young (≤ 10^{3} yr).”

SUPERNOVA 1995E IN NGC 2441 P. Garnavich, A. Riess, P. Challis, and R. Kirshner, Center for Astro-

physics, report that this object is highly reddened. A spectrum obtained by P. Berlind on Feb. 24 UT with the 1.5-m Tillinghast Telescope on Mt. Hopkins shows that SN 1995E has an exceptionally red continuum for a normal type-Ia supernova near maximum. In addition, a strong, narrow absorption feature due to Na I is seen near the redshift of NGC 2441. The measured equivalent width of 0.14 nm, combined with the continuum shape, suggests that the visual extinction to SN 1995E could well exceed one magnitude.

NOTICE TO CONTRIBUTORS A planned electrical shutdown means that the Central Bureau will be

unable to react quickly to electronic communications on Mar. 4 (Saturday), possibly for the whole of the UT day. Incoming messages will be stored for later attention, and it is probable that the shutdown will not in fact affect the whole day. Telephones will be manned for some of the downtime, so really urgent information can be reported in this manner.

1995 March 2 Daniel W. E. Green